The height of a cylinder is twice the diameter. Express the total surface area A as a function of the height h.
This question really stump me can you show the work as well for me?

Answer by scott8148(6628)

"The height of a cylinder is twice the diameter" ___ h = 2d ___ h/2 = d ___ h/4 = r

A = LA + ends = [2 * pi * r * h] + [2 * pi * r^2] = [2 * pi * h/4 * h] + [2 * pi * (h/4)^2]

A = [2 * pi * (h^2 / 4)] (5 / 4) = 5/8 (pi * h^2)